OpenStudy (tmagloire1):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(2t). Find the acceleration of the particle when the velocity is first zero. 2e^2 2e e None of these

Vocaloid (vocaloid):

well, we can find the derivative of position to get velocity, then we can take the derivative of velocity to get acceleration do you know how to find the derivative of t*ln(2t)? we'll be using the product rule

OpenStudy (tmagloire1):

@Vocaloid log(2t)+1

Vocaloid (vocaloid):

almost, ln(2t) + 1

OpenStudy (tmagloire1):

oops i meant ln xD

Vocaloid (vocaloid):

no prob, now all we need to do is take the derivative of ln(2t). can you do that?

OpenStudy (tmagloire1):

So the next derivative is 1/t

Vocaloid (vocaloid):

yup! our last step is to figure out what to plug in for t the problem wants to find the acceleration when the velocity is 0, so we go back to our velocity equation and set it equal to 0, then solve for t ln(2t) + 1 = 0 solve for t

OpenStudy (tmagloire1):

i got 1/2e

Vocaloid (vocaloid):

right, t = 1/(2e), now we plug that back into our acceleration equation 1/t = 1/(1/(2e) = 1/(0.5e^-1) = 2e

OpenStudy (tmagloire1):

Thank you so much for the help!